### v1

# Weighted Q-learning for optimal dynamic treatment regimes with MNAR coavriates

Abstract
Dynamic treatment regimes (DTRs) formalize medical decision-making as a sequence of rules for different stages, mapping patient-level information to recommended treatments. In practice, estimating an optimal DTR using observational data from electronic medical record (EMR) databases can be complicated by covariates that are missing not at random (MNAR) due to informative monitoring of patients. Since complete case analysis can result in consistent estimation of outcome model parameters under the assumption of outcome-independent missingness (Yang et al., 2019), Q-learning is a natural approach to accommodating MNAR covariates. However, the backward induction algorithm used in Q-learning can introduce complications, as MNAR covariates at later stages can result in MNAR pseudo-outcomes at earlier stages, leading to suboptimal DTRs, even if outcome variables are fully observed. To address this unique missing data problem in DTR settings, we propose two weighted Q-learning approaches where inverse probability weights for missingness of the pseudo-outcomes are obtained through estimating equations with valid nonresponse instrumental variables or sensitivity analysis. Asymptotic properties of the weighted Q-learning estimators are derived and the finite-sample performance of the proposed methods is evaluated and compared with alternative methods through extensive simulation studies. Using EMR data from the Medical Information Mart for Intensive Care database, we apply the proposed methods to investigate the optimal fluid strategy for sepsis patients in intensive care units.
Keywords Backward-induction induced missing pseudo-outcome; Future-independent missingness; Missing not at random; Nonresponse instrumental variable; Q-learning; Sensitivity analysis.
1 Introduction
Dynamic treatment regimes (DTRs) formalize medical decision-making as a sequence of decision rules, each corresponding to a specific decision stage, that map available information at patient-level to a recommended treatment (Murphy, 2003). DTRs have been estimated and evaluated in various biomedical areas, including cancer (Xu et al., 2019), HIV (Yu et al., 2019), and sepsis (Yu and Zhang, 2022), to name just a few. In practice, estimating DTRs becomes challenging when there are missing data in covariates, treatments, or outcomes.
There is a rich literature on identifying optimal DTRs with complete data. Various methods have been developed, including Q-learning (Murphy, 2003; Nahum-Shani et al., 2012), A-learning (Schulte et al., 2014), and their numerous variants (Goldberg and Kosorok, 2012; Shi et al., 2018; Kosorok and Laber, 2019). These methods rely on approximate dynamic programming methods. Direct-optimization methods, such as simultaneous outcome weighted learning (Zhao et al., 2015), and model-based planning via g-computation (Xu et al., 2016; Laber et al., 2018) have also been employed. In contrast, limited research to date has been devoted to accommodating missing data when estimating DTRs. For survival outcomes, a few methods have been proposed to estimate DTRs by addressing missingness due to right-censoring (Simoneau et al., 2020; Cho et al., 2023). For handling missing data in covariates, treatments and outcomes in general, Shortreed et al. (2014) proposed a time-ordered nested conditional imputation strategy when estimating DTRs in sequential multiple assignment randomized trials. Dong et al. (2020) applied the augmented inverse probability weighting method to Q-learning and a generalized version of outcome-weighted learning when dealing with missing data due to patients’ dropout. In simpler settings of estimating single-stage optimal treatment regimes, Shen et al. (2023) further discussed the multiple imputation method proposed by Shortreed et al. (2014) for estimating and evaluating individualized treatment rules that were not directly observed in the design. Also for the single-stage setting, Huang and Zhou (2020) investigated the performance of an augment inverse probability weighted estimator in the direct-optimization framework when covariates are missing at random. However, these existing methods all focus on data missing at random and could potentially yield sub-optimal DTRs when data for covariates, treatments or outcomes are missing not at random (MNAR) (Little and Rubin, 2014).
In this research, we aim to address the issue of MNAR covariates in the estimation of DTRs, which is prominent in settings where patients are informatively monitored, e.g., in electronic medical records (EMR) databases or disease clinics. Our research was motivated by a recent investigation on optimal fluid strategy for sepsis patients in intensive care units (ICUs) (Speth et al., 2022). Using EMR data from the Medical Information Mart for Intensive Care (MIMIC-III) (Johnson et al., 2016), Speth et al. (2022) estimated an optimal two-stage DTR for treating sepsis patients with fluid resuscitation in ICUs by incorporating baseline covariates such as demographics, time-varying covariates such as mechanical ventilation usage and vasopressor administration, and the outcome measure of multi-organ failure at the first stage. However, according to the recommendations of the ‘Surviving Sepsis Campaign’, hemodynamic variables should also be taken into account when reassessing the need for additional fluid administration following the initial resuscitation (Rhodes et al., 2017). Unfortunately, within the MIMIC-III database, certain hemodynamic variables, including blood pressure, respiratory rate, and body temperature, are subject to missingness at the conclusion of the initial resuscitation. Furthermore, as highlighted by Awad et al. (2017), patients in the MIMIC-III database were informatively monitored since patients with more severe conditions, who typically presented with more abnormal physiological indicators than others, received more intensive monitoring. Consequently, their measurements of hemodynamic variables were less likely to be missing. In this case, the missingness probabilities of the hemodynamic variables are directly related to their own values, indicating that the missing data mechanism is MNAR. This characteristic could potentially lead to sub-optimal estimation of DTRs if existing methods are employed.
The identification and estimation of DTRs become much more challenging when covariates are MNAR. In the literature on handling MNAR covariates, it has been concluded that the parameters of an outcome regression model can be consistently estimated using complete case analysis as long as the missingness of the covariates is not directly related to the outcome (Yang et al., 2019). Therefore, considering that Q-learning relies on consistently estimated outcome models and is both flexible and straightforward to implement, it offers a natural approach for accommodating MNAR covariates when estimating DTRs. However, when implementing the step-wise outcome regression algorithm in Q-learning, challenges arise not only from the missingness of the covariates but also from the missingness of the pseudo-outcomes, even if outcome variables are fully observed. This is because the Q-learning algorithm computes the pseudo-outcome in the preceding stage using covariates from the subsequent stage. As a result, if there are MNAR covariates in the later stages, it can lead to MNAR pseudo-outcomes in the earlier stages. We provide a detailed illustration of this unique missing data problem with DTRs caused by backward-induction-induced MNAR pseudo-outcomes in Section 2.3.
It is well known that with MNAR outcomes, outcome regression models are generally not identifiable; see examples provided by Miao et al. (2016). Inspired by the MNAR outcome literature, we propose two weighted Q-learning approaches to tackling MNAR pseudo-outcomes in DTR estimation. The first approach is to employ estimating equations to estimate a working model for missingness probability of the pseudo-outcome with valid nonresponse instrumental variables (D’Haultfoeuille, 2010; Wang et al., 2014; Miao and Tchetgen Tchetgen, 2016). The pseudo-outcome estimates based on observed data are then weighted by the inverse of the estimated missingness probabilities in the Q-learning algorithm. In scenarios where nonresponse instrumental variables are not available, we specify a sensitivity parameter to quantify the residual association between the missing pseudo-outcome at stage $t$ and its conditional missingness probability given observed information up to stage $t$. Thus the missingness probabilities of the pseudo-outcomes are functions of the sensitivity parameter. We then perform a sensitivity analysis to evaluate the robustness of the estimated optimal DTRs with respect to the pre-specified sensitivity parameter (Rotnitzky et al., 1998; Kim and Yu, 2011; Wen and Seaman, 2018).
2 Method
2.1 Setting and notation
We consider an observational cohort with $n$ patients and finite $T$ treatment stages. Let subscripts $i=1,\ldots,n$, $t=1,\ldots,T$ denote patients and stages, respectively. Unless specified otherwise, we use capital letters to denote random variables and lowercase letters to indicate specific realizations of random variables. Complete data for the patients are assumed independent and identically distributed and thus we suppress the patient-specific subscript $i$. Let $A_{t}\in\{-1,1\}$ be the assigned treatment at stage $t$ for $t=1,\ldots,T$. $\bm{X}_{t}\in\mathbb{R}^{p_{t}}$ is a $1\times p_{t}$ vector of covariates measured before $A_{t}$, and $Y_{t}$ denotes the longitudinal outcome measured after $A_{t}$ at stage $t$.
The final outcome of interest is a pre-specified function (e.g., sum or maximum) of the longitudinal outcome in $T$ stages, $Y=f(Y_{1},\ldots,Y_{T})$. Higher values of $Y$ are assumed to be better. We define $\bm{H}_{1}=\bm{X}_{1}$, and $\bm{H}_{t}=\left(\bm{H}_{t-1},A_{t-1},Y_{t-1},\bm{X}_{t}\right)$ for $t=2,\ldots,T$. Thus, $\bm{H}_{t}$ represents the information available before making the treatment decision at stage $t$. Let $\mathcal{H}_{t}$ denote the support of $\bm{H}_{t}$. A DTR $\bm{d}$ consists of a set of decision rules $(d_{1},\ldots,d_{T})$, where $d_{t}:\mathcal{H}_{t}\to\{-1,1\}$ is a function that takes the observed history $\bm{h}_{t}$ as input and outputs a treatment decision at stage $t$. An optimal DTR is the set of decision rules $\left(d^{\mathrm{opt}}_{1},\ldots,d^{\mathrm{opt}}_{T}\right)$ that maximizes the expectation of the final outcome. We formalize this definition using the potential outcome framework.
For $t=1,\ldots,T$, let $Y_{t}^{*}(\bar{\bm{a}}_{t})$ denote the potential longitudinal outcome at stage $t$ and $\bm{X}_{t+1}^{*}(\bar{\bm{a}}_{t})$ denote the potential covariates at stage $t+1$ if a patient, possibly contrary to fact, had received the treatment sequence $\bar{\bm{a}}_{t}=(a_{1},\ldots,a_{t})$ by stage $t$. We then define the set of potential outcomes under $\bar{\bm{a}}_{t}$ as $\bm{O}_{t}^{*}(\bar{\bm{a}}_{t})=\{Y_{1}^{*}(a_{1}),\bm{X}_{2}^{*}(a_{1}),\ldots,Y_{t}^{*}(\bar{\bm{a}}_{t}),\bm{X}_{t+1}^{*}(\bar{\bm{a}}_{t})\}$ for $t=1,\ldots,T-1$. The potential final outcome under a regime $\bm{d}$ is
$Y^{*}(\bm{d})=\sum_{\bar{\bm{a}}_{T}}Y^{*}\left(\bar{\bm{a}}_{T}\right)\prod_{t=1}^{T}\mathbb{I}[d_{t}\left\{(\bm{X}_{1},a_{1},Y_{1}^{*}(a_{1}),\ldots,a_{t-1},Y_{t-1}^{*}(\bar{\bm{a}}_{t-1}),\bm{X}_{t}^{*}(\bar{\bm{a}}_{t-1}))\right\}=a_{t}]$
where $Y^{*}\left(\bar{\bm{a}}_{T}\right)=f\left(Y_{1}^{*}\left(a_{1}\right),Y_{2}^{*}\left(\bar{\bm{a}}_{2}\right),\ldots,Y_{T}^{*}\left(\bar{\bm{a}}_{T}\right)\right)$. Define the value of a regime $\bm{d}$ to be $V(\bm{d})=E[Y^{*}(\bm{d})]$. Let $\mathcal{D}$ denote all feasible regimes. An optimal DTR, $\bm{d}^{\mathrm{opt}}\in\mathcal{D}$, satisfies that $V(\bm{d}^{\mathrm{opt}})\geq V(\bm{d})$ for all $\bm{d}\in\mathcal{D}$.
2.2 Q-learning with complete data
In this section, we briefly review the Q-learning algorithm for estimating optimal DTRs with complete observational data. To estimate the value of candidate regimes from observed data, it is necessary to express the value of regimes solely in terms of observables rather than potential outcomes. This becomes feasible under the following, now standard, causal inference assumptions:
Assumption 1 (Consistency). $Y_{t}^{*}(\bar{\bm{a}}_{t})=Y_{t}$ for $t=1,\ldots,T$ and $\bm{X}_{t+1}^{*}(\bar{\bm{a}}_{t})=\bm{X}_{t+1}$ for $t=1,\ldots,T-1$ when $\bar{\bm{a}}_{t}$ are actually received.
[…]
Assumption 3 (Positivity). $P(A_{t}=a_{t}\mid\bm{H}_{t})>0$ for $a_{t}\in\{-1,1\}$, $\bm{H}_{t}\in\mathcal{H}_{t}$ and $t=1,\ldots,T$.
Under Assumptions 1-3, the value of a DTR can be identified from the observed data. In particular, for the 2-stage scenario we have that
$E[Y^{*}(\bm{d})]=E[E\{E(Y\mid\bm{H}_{2},A_{2}=d_{2}(\bm{H}_{2}))\mid\bm{H}_{1},A_{1}=d_{1}(\bm{H}_{1})\}]$
This repeated expectation form suggests that the optimal DTR can be computed with a backward-induction procedure, which is expressed in terms of the following recursively defined Q-functions in canonical Q-learning:
$Q_{T}\left(\bm{h}_{T},a_{T}\right)$ $=E\left(Y\mid\bm{H}_{T}=\bm{h}_{T},A_{T}=a_{T}\right),$
$Q_{t}\left(\bm{h}_{t},a_{t}\right)$ $=E\left\{\max_{a_{t+1}\in\{-1,1\}}Q_{t+1}\left(\bm{H}_{t+1},a_{t+1}\right)\mid\bm{H}_{t}=\bm{h}_{t},A_{t}=a_{t}\right\},$
for $t=T-1,\ldots,1$.
The true Q-functions are unknown and must be estimated from the data. Since Q-functions represent conditional expectations, it is natural to estimate them using regression models. Note that in these models, the response variable is $\max_{a_{t+1}\in\{-1,1\}}Q_{t+1}\left(\bm{H}_{t+1},a_{t+1}\right)$ instead of $Y_{t}$ at stage $t=1,\ldots,T-1$. These response variables are commonly referred to as pseudo-outcomes, and we define $Y_{pse,T}=Y$ and $Y_{pse,t}=\max_{a_{t+1}\in\{-1,1\}}Q_{t+1}\left(\bm{H}_{t+1},a_{t+1}\right)$ for $t=1,\ldots,T-1$. Subsequently, $Q_{t}(\bm{h}_{t},a_{t})=E\left(Y_{pse,t}\mid\bm{h}_{t},a_{t}\right)$ for $t=1,\ldots,T$. Since $a_{t}\in\{-1,1\}$ , for any specified model for $Q_{t}(\bm{h}_{t},a_{t})$, we can decompose it as $Q_{t}(\bm{h}_{t},a_{t};\bm{\theta}_{t})=q_{t,0}(\bm{h}_{t};\bm{\beta}_{t})+a_{t}q_{t,1}(\bm{h}_{t};\bm{\psi}_{t})$, where $\bm{\theta}_{t}=(\bm{\beta}_{t},\bm{\psi}_{t})$, $q_{t,0}(\cdot)=\frac{Q_{t}(\bm{h}_{t},1)+Q_{t}(\bm{h}_{t},-1)}{2}$, and $q_{t,1}(\cdot)=\frac{Q_{t}(\bm{h}_{t},1)-Q_{t}(\bm{h}_{t},-1)}{2}$ for $t=1\ldots,T$. As a result, the Q-function is divided into two components: (1) a stage $t$ treatment-free component $q_{t,0}(\bm{h}_{t};\bm{\beta}_{t})$ which depends on (a subset of) the history $\bm{h}_{t}$ before stage $t$ but not on the stage $t$ treatment $a_{t}$, and (2) a stage $t$ treatment effect component $a_{t}q_{t,1}(\bm{h}_{t};\bm{\psi}_{t})$ which depends on (a potentially different subset of) $\bm{h}_{t}$ and specifically includes the main effect of treatment $a_{t}$ and its interactions with tailoring variables (i.e., effect modifiers). The blip function, as defined in Chakraborty and Moodie (2013), equals to $a_{t}q_{t,1}(\bm{h}_{t};\bm{\psi}_{t})+q_{t,1}(\bm{h}_{t};\bm{\psi}_{t})$ under this setting. Therefore, we refer to $\bm{\psi}_{t}$ as the blip function parameters. Using this modelling approach, we have $d_{t}^{\mathrm{opt}}(\bm{h}_{t})=2\mathbb{I}\left(q_{t,1}(\bm{h}_{t};\bm{\psi}_{t})>0\right)-1$. Under Assumptions 1-3, the estimator $\hat{\bm{\theta}}_{t}$ of $\bm{\theta}_{t}$ can be recursively constructed by regressing $Y_{pse,t}$ on $\bm{H}_{t}$ and $A_{t}$ for $t=T,\ldots,1$. Then, the Q-learning estimator of $d^{\mathrm{opt}}_{t}$ can be obtained by $\hat{d}_{t}^{\mathrm{opt}}\left(\bm{h}_{t}\right)=2\mathbb{I}\left(q_{t,1}(\bm{h}_{t};\hat{\bm{\psi}}_{t})>0\right)-1$ for $t=1,\ldots,T$.
2.3 Q-learning for DTRs with MNAR covariates: why complete case analysis fails at earlier stages while working at the last stage?
We focus on the scenario that the treatment and longitudinal outcome, $(A_{t},Y_{t})$, are fully observed while the covariates, $\bm{X}_{t}$, contain missing values. This scenario is frequently encountered in practice because treatments are typically well-documented and the longitudinal outcomes of interest are usually key clinical disease activity indices which are regularly measured. Moreover, the probability of a covariate being missing is likely to depend on the value of the covariate itself due to informative monitoring of patients. For instance, in the MIMIC-III data, patients who displayed symptoms of abnormal heart rates such as chest pain were more likely to have their heart rates recorded. Therefore, the lack of heart rate measurements could indicate MNAR, since patients with normal heart rates were less likely to be frequently monitored.
Let $R_{t}$ represent the missingness indicator for $\bm{X}_{t}$ such that $R_{t}=1$ if all variables in $\bm{X}_{t}$ are fully observed and $R_{t}=0$ if at least one component of $\bm{X}_{t}$ is missing. Define $\bar{\bm{R}}_{t}=(R_{1},\ldots,R_{t})$ as the history of missingness indicators up to stage $t$. We use $\bm{1}_{t}$ to denote a $t$-dimensional vector in which all elements are equal to $1$ for $t=1,\ldots,T$. In ICU settings, the effect of a treatment is commonly evaluated by the patient’s longitudinal outcomes (e.g., physiological symptoms) observed several hours after treatment administration, while the treatment decision for a patient is based on the patient’s medical history. Therefore, it is plausible that the missing data mechanism of the medical history $\bm{H}_{t}$ does not directly depend on the patient’s variables measured after receiving the current stage treatment $A_{t}$. Thus we make the following assumption:
[…]
Figure 1 is a directed acyclic graph illustrating Assumptions 2 and 4 in a 2-stage scenario. In this example, $\bm{H}_{1}=\bm{X}_{1}$, $\bm{H}_{2}=(\bm{X}_{1},A_{1},Y_{1},\bm{X}_{2})$. $R_{1}$ and $R_{2}$ are missing indicators for $\bm{X}_{1}$ and $\bm{X}_{1}$, respectively. Graphically, $A_{1}$ has no common parents with ($Y_{1},\bm{X}_{2},A_{2},Y_{2}$) except for $\bm{H}_{1}$, and $A_{2}$ has no common parents with $Y_{2}$ except for $\bm{H}_{2}$, encoding Assumption 2. Besides, $R_{1}$ is d-separated with $(Y_{1},\bm{X}_{2},R_{2},A_{2},Y_{2})$ by $(\bm{H}_{1},A_{1})$, and $(R_{1},R_{2})$ is d-separated with $Y_{2}$ by $(\bm{H_{2}},A_{2})$ (VanderWeele and Robins, 2007), encoding Assumption 4.
Figure 1: A direct acyclic graph illustrating Assumptions 2 and 4
[…]
However, conditional on $\bar{\bm{R}}_{t}=\bm{1}_{t}$ (i.e., for complete cases up to stage $t$), although there is no missingness in $\bm{H}_{t}$, the pseudo-outcome $Y_{pse,t}$ may contain missing values because $Y_{pse,t}$ depends on $\bm{X}_{t+1}$, which can be missing (i.e., $R_{t+1}=0$) when $\bar{\bm{R}}_{t}=\bm{1}_{t}$. We illustrate this backward-induction induced missing pseudo-outcome problem in Figure 2 with a hypothetical example coming from the same setting of our simulation study in Section 3. In this example, all variables for Patient 3 at stage one are fully observed. Nevertheless, $Y_{pse,1}$ for Patient 3 is unavailable because $X_{22}$, a time-varying covariate required for computing $Y_{pse,1}$, is missing.
Figure 2: A hypothetical example illustrating the relationship between the missing pattern of covariates and that of pseudo-outcomes. Note: Gray areas represent missing information. The setting of the variables is the same as that in Simulation One in Section 3.
Let $R_{pse,t}$ denote the missingness indicator of $Y_{pse,t}$ for $t=1,\ldots,T$. Because the missingness probabilities of $\bm{X}_{t+1}$ depend on their own values, and $Y_{pse,t}$ is computed based on $\bm{X}_{t+1}$, it is likely that $Y_{pse,t}$ is missing not at random. This backward-induction-induced MNAR pseudo-outcome problem would make identification and consistent estimation of Q-functions impossible. A complete case analysis for Q-learning would likely result in suboptimal DTR. In Section 3, we will use simulations to demonstrate this phenomenon.
2.4 Weighted Q-learning for DTR with MNAR pseudo-outcomes: a nonresponse instrument variable approach
We propose two approaches for inverse probability weighted Q-learning to tackle the MNAR pseudo-outcome problem. The first is based on valid nonresponse instrumental variables. Specifically, we consider the following semiparametric model for $R_{pse,t}$:
Assumption 5 (semiparametric working model for missingness probability of the pseudo-outcomes). For $t=1,\ldots,T-1$,
$P(R_{pse,t}=1\mid\bm{h}_{t},a_{t},y_{pse,t},\bar{\bm{R}}_{t}=\bm{1}_{t})=\pi(\bm{u}_{t},y_{pse,t};\gamma_{t})=\frac{1}{1+\exp\{s_{t}(\bm{u}_{t})+\gamma_{t}y_{pse,t}\}},$ (1)
where $\bm{U}_{t}\subset(\bm{H}_{t},A_{t})$ is a proper subset of $(\bm{H}_{t},A_{t})$, $P(Y_{pse,t}\mid\bm{H}_{t},A_{t})\neq P(Y_{pse,t}\mid\bm{U}_{t})$, and $s_{t}(\cdot)$ is an unknown and unspecified function of $\bm{u}_{t}$.
This semiparametric model assumption does not impose any parametric constraint on the function of $\bm{u}_{t}$ in working model (1). Instead, it restricts the parametric form of $y_{pse,t}$ and requires that there is no interaction between $\bm{u}_{t}$ and $y_{pse,t}$ in working model (1). It is noteworthy that the variables observed after stage $t$ may exhibit an association with $R_{pse,t}$ even after conditioning on ($\bm{h}_{t},a_{t},y_{pse,t},\bar{\bm{R}}_{t}=\bm{1}_{t}$) and therefore model (1) may not be the true data-generating model for $R_{pse,t}$. However, weighting by the inverse of $P(R_{pse,t}=1\mid\bm{h}_{t},a_{t},y_{pse,t},\bar{\bm{R}}_{t}=\bm{1}_{t})$ would suffice to balance the distribution of the variables between the patients with and without missing $y_{pse,t}$ for the purpose of estimating $Q_{t}(\bm{h}_{t},a_{t})$ and thus lead to unbiased estimates of the Q-functions after weighting.
[…]
If Assumptions 5 holds, to estimate the working model (1), we have to overcome the challenge that $s_{t}(\cdot)$ is nonparametric. Following Shao and Wang (2016), we consider the idea of profiling and obtain the following kernel estimate of $s_{t}(\cdot)$ for a given $\gamma_{t}$.
$\exp\{\hat{s}_{\gamma_{t},t}(\bm{u}_{t})\}=\frac{\sum_{i=1}^{n}\left(1-r_{pse,t,i}\right)K_{c_{t}}\left(\bm{u}_{t}-\bm{u}_{t,i}\right)\mathbb{I}(\bar{\bm{R}}_{t,i}=\bm{1}_{t})}{\sum_{i=1}^{n}r_{pse,t,i}\exp\left(\gamma_{t}y_{pse,t,i}\right)K_{c_{t}}\left(\bm{u}_{t}-\bm{u}_{t,i}\right)\mathbb{I}(\bar{\bm{R}}_{t,i}=\bm{1}_{t})},$ (2)
where $K_{c_{t}}(\cdot)=c_{t}^{-1}K(\cdot/c_{t})$, with $K(\cdot)$ being a symmetric kernel function and $c_{t}$ a bandwidth. For a fixed $\gamma_{t}$, the working model (1) can be expressed as $\hat{\pi}_{t}(\bm{u}_{t},y_{pse,t};\gamma_{t})=[1+\exp\{\hat{s}_{\gamma_{t},t}(\bm{u}_{t})+\gamma_{t}y_{pse,t}\}]^{-1}$. Then, we can estimate $\gamma_{t}$ with the following estimating equation,
$\widehat{E}\left[\mathbb{I}(\bar{\bm{R}}_{t}=\bm{1}_{t})\bm{l}_{t}(\bm{z}_{t})\left\{\frac{r_{pse,t}}{\hat{\pi}_{t}(\bm{u}_{t},y_{pse,t};\gamma_{t})}-1\right\}\right]=\bm{0},$ (3)
where $\widehat{E}[\cdot]$ denotes the empirical mean, $\bm{l}_{t}(\bm{z}_{t})=(l_{t,1}(\bm{z}_{t}),\ldots,l_{t,L}(\bm{z}_{t}))$ is a user-specified differentiable vector function of the nonresponse instrument variables of dimension $L$, $L\geq 2$.
Note that in estimating equations (3), if $\bm{z}_{t}$ are constants, then the above equations are not solvable because they are under-identified. Nevertheless, when $L>2$ and the equations are over-identified, we can employ the two-step generalized method of moments (GMM) (Hansen, 1982) to obtain the estimator of $\gamma_{t}$. Let $b_{t,j}(r_{pse,t},\bm{u}_{t},y_{pse,t};\gamma_{t})=\mathbb{I}(\bar{\bm{R}}_{t}=\bm{1}_{t})l_{t,j}(\bm{z}_{t})\left(\frac{r_{pse,t}}{\hat{\pi}_{t}(\bm{u}_{t},y_{pse,t};\gamma_{t})}-1\right)$ for $j=1,\ldots,L$ and $\bm{B}_{t}(\gamma_{t})=(\widehat{E}[b_{t,1}],\widehat{E}[b_{t,2}],\ldots,\widehat{E}[b_{t,L}])$. Then, the first-step GMM estimator of $\gamma_{t}$ is $\hat{\gamma}^{(1)}_{t}=\underset{\gamma_{t}}{\arg\min}\;\bm{B}_{t}(\gamma_{t})\bm{B}_{t}(\gamma_{t})^{\mathrm{T}}$.
Based on the first-step GMM estimator, we can build a more efficient estimator of $\gamma_{t}$ with a weight matrix $\bm{W}_{t}$. Let $\bm{W}_{t}$ be the inverse matrix of the $L\times L$ matrix with the $\left(j,j^{\prime}\right)$ element as $\frac{1}{n}\sum_{i=1}^{n}b_{t,j}(r_{pse,t},\bm{u}_{t},y_{pse,t};\hat{\gamma}^{(1)}_{t})b_{t,j^{\prime}}(r_{pse,t},\bm{u}_{t},y_{pse,t};\hat{\gamma}^{(1)}_{t})$. Then the second-step GMM estimator of $\gamma_{t}$ is $\hat{\gamma}_{t}=\underset{\gamma_{t}}{\arg\min}\;\bm{B}_{t}(\gamma_{t})\bm{W}_{t}\bm{B}_{t}(\gamma_{t})^{\mathrm{T}}$.
With the estimated working model for the missingness probability, we can employ the weighted least square method to obtain unbiased estimates of the Q-function parameters. Subsequently, the optimal treatment at stage $t$ can be determined based on the estimated parameters. We henceforth refer to this approach as the estimating equation (EE) method.
2.5 Weighted Q-learning for DTRs with MNAR pseudo-outcomes: a sensitivity analysis approach
In real data applications, it might be challenging to find valid nonresponse instrument variables for every stage based on domain knowledge. In such scenarios, we develop a sensitivity analysis approach for the missingness probability of $Y_{pse,t}$. Specifically, we slightly modify Assumption 5 and introduce the following assumption.
Assumption 6 (Semiparametric working model for missingness probability without nonresponse instrumental variables).
$P(R_{pse,t}=1\mid\bm{h}_{t},a_{t},y_{pse,t},\bar{\bm{R}}_{t}=\bm{1}_{t})=\pi(\bm{h}_{t},a_{t},y_{pse,t};\gamma_{t}^{\prime})=\frac{1}{1+\exp\{s^{\prime}_{t}(\bm{h}_{t},a_{t})+\gamma^{\prime}_{t}y_{pse,t}\}},$ (4)
where $s^{\prime}_{t}(\cdot)$ is an unknown and unspecified function of $(\bm{h}_{t},a_{t})$ and $\gamma^{\prime}_{t}$ is the pre-specified sensitivity parameter. Note that here we include all historical variables in (4).
The sensitivity parameter $\gamma^{\prime}_{t}$ quantifies the residual effect of $Y_{pse,t}$ on $R_{pse,t}$ after adjusting for $\bm{h}_{t}$ and $a_{t}$ for patients with complete records up to stage $t$. The working model (4) can be identified with a fixed $\gamma^{\prime}_{t}$ (Kim and Yu, 2011). We can obtain the kernel estimate of $\exp\{s^{\prime}_{\gamma^{\prime}_{t},t}(\bm{h}_{t})\}$ in a similar manner as in equation (2).
$\exp\{\hat{s}^{\prime}_{\gamma^{\prime}_{t},t}(\bm{h}_{t})\}=\frac{\sum_{i=1}^{n}\left(1-r_{pse,t,i}\right)K_{c^{\prime}_{t}}\left(\bm{h}_{t}-\bm{h}_{t,i}\right)\mathbb{I}(\bar{\bm{R}}_{t,i}=\bm{1}_{t})}{\sum_{i=1}^{n}r_{pse,t,i}\exp\left(\gamma^{\prime}_{t}y_{pse,t,i}\right)K_{c^{\prime}_{t}}\left(\bm{h}_{t}-\bm{h}_{t,i}\right)\mathbb{I}(\bar{\bm{R}}_{t,i}=\bm{1}_{t})},$ (5)
where $K_{{c^{\prime}}_{t}}(\cdot)={c^{\prime}}_{t}^{-1}K(\cdot/c^{\prime}_{t})$, with $K(\cdot)$ being a symmetric kernel function and $c^{\prime}_{t}$ a bandwidth.
Subsequently, we can employ the inverse probability weighting method to obtain unbiased estimates of the Q-functions and the optimal DTRs. If the working model (4) is correctly specified (including the sensitivity parameter $\gamma^{\prime}_{t}$) and certain regularity conditions for $\exp\{s^{\prime}_{\gamma^{\prime}_{t},t}(\bm{h}_{t})\}$ are satisfied, along with correctly specified Q-functions, the parameter estimators in the Q-functions are consistent. As the true value of $\gamma^{\prime}_{t}$ is unknown, we conduct a sensitivity analysis by varying the value of $\gamma^{\prime}_{t}$ to assess the robustness of the Q-learning results with respect to $\gamma^{\prime}_{t}$. We refer to this approach as the SA method.
2.6 Summary of the estimation steps
We summarize the estimation step of the proposed weighted Q-learning approach as follows:
  1. 1.
    Compute $y_{pse,T}=f(y_{1},y_{2},\ldots,y_{T})$. Let $\bm{M}_{T}(\bm{\theta}_{T})=\widehat{E}\left\{(\bar{\bm{R}}_{T}=\bm{1}_{T})\frac{\partial\left[\left\{Q_{T}(\bm{h}_{T},a_{T};\bm{\theta}_{T})-y_{pse,T}\right\}^{2}\right]}{\partial\bm{\theta}_{T}}\right\}$. Then obtain $\hat{\bm{\theta}}_{T}$ by solving the estimating equation $\bm{M}_{T}(\bm{\theta}_{T})=\bm{0}$
  2. 2.
    For $t=T-1,\ldots,1$, repeat the following steps recursively:
    1. (a)
      Construct stage $t$ pseudo-outcome as $\hat{y}_{pse,t}=\max\limits_{a_{t+1}}\;Q_{t+1}(\bm{h}_{t+1},a_{t+1};\hat{\bm{\theta}}_{t+1})$, denote the missingness indicator of $\hat{y}_{pse,t}$ as $\hat{r}_{pse,t}$.
    2. (b)
      Estimate $P(\hat{r}_{pse,t}\mid\bm{h}_{t},a_{t},\hat{y}_{pse,t},\bar{\bm{R}}_{t}=\bm{1}_{t})$ by (i) solving semi-parametric estimating equations (3) with the EE method under Assumption 5 or (ii) the SA method under Assumption 6 at stage $t$. Denote this estimated conditional probability as $\hat{\pi}_{t}$.
    3. (c)
      Let $\bm{M}_{t}(\bm{\theta}_{t})=\widehat{E}\left\{\frac{\hat{r}_{pse,t}}{\hat{\pi}_{t}}\mathbb{I}(\bar{\bm{R}}_{t}=\bm{1}_{t})\frac{\partial\left[\left\{Q_{t}(\bm{h}_{t},a_{t};\bm{\theta}_{t})-\hat{y}_{pse,t}\right\}^{2}\right]}{\partial\bm{\theta}_{t}}\right\}$. Solve the estimating equation
      $\bm{M}_{t}(\bm{\theta}_{t})=\bm{0}$ to obtain $\hat{\bm{\theta}}_{t}$.
  3. 3.
    The optimal treatment regime at stage $t$ can be obtained by $\hat{d}_{t}^{\mathrm{opt}}\left(\bm{h}_{t}\right)=\operatorname*{arg\,max}\limits_{a_{t}}Q_{t}(\bm{h}_{t},a_{t};\hat{\bm{\theta}}_{t})$ for $t=1,\ldots,T$.
Remark 1. Recall that, since $a_{t}\in\{-1,1\}$, $Q_{t}(\bm{h}_{t},a_{t};\hat{\bm{\theta}}_{t})$ can be written as $q_{t,0}(\bm{h}_{t};\hat{\bm{\beta}}_{t})+a_{t}q_{t,1}(\bm{h}_{t};\hat{\bm{\psi}}_{t})$. Subsequently, $\hat{y}_{pse,t}$ can be calculated by $q_{t+1,0}(\bm{h}_{t+1};\hat{\bm{\beta}}_{t+1})+|q_{t+1,1}(\bm{h}_{t+1};\hat{\bm{\psi}}_{t+1})|$ and $\hat{d}_{t}^{\mathrm{opt}}\left(\bm{h}_{t}\right)$ is determined by the sign of $q_{t,1}(\bm{h}_{t};\hat{\bm{\psi}}_{t})$.
Remark 2. The pseudo-outcome $Y_{pse,t}$ is determined by the specified parametric form of $Q_{t+1}(\cdot)$ and the true values of $\bm{\theta}_{t+1}$ for $t=1,\ldots,T-1$. Since the true value of $\bm{\theta}_{t+1}$ is unknown in practice, we use $\hat{Y}_{pse,t}$ calculated by the parameter estimates $\hat{\bm{\theta}}_{t+1}$ in the estimation steps. In this scenario, with the EE method, $P(\hat{Y}_{pse,t}\mid\bm{H}_{t},A_{t})$ and $\pi(\bm{u}_{t},\hat{y}_{pse,t})$ might be slightly different from $P(Y_{pse,t}\mid\bm{H}_{t},A_{t})$ and $\pi(\bm{u}_{t},y_{pse,t})$, respectively. Nevertheless, if all the models are correctly specified, $\hat{\bm{\theta}}_{t+1}$ serves as a consistent estimator for $\bm{\theta}_{t+1}$. Subsequently, $\hat{Y}_{pse,t}$ converges to $Y_{pse,t}$. It follows that the estimated optimal DTR $\hat{\bm{d}}$ will also converge to the true optimal DTR $\bm{d}$. We formalize these asymptotic properties in the following section. In Section 3, we also conduct simulations to assess the robustness of the proposed EE method against mild model misspecification, e.g., when the chosen instrumental variable is weakly associated with $R_{pse,t}$ after conditioning on $\bm{U}_{t}$ and $Y_{pse,t}$.
We provide an R Markdown tutorial for implementing the weighted Q-learning with the EE method using simulated data in the Supplementary Materials.
2.7 Asymptotic properties
The consistency and asymptotic normality of the estimators in Section 2.6 can be established under suitable regularity conditions for estimating equations and an additional assumption that the optimal treatment is unique for all subjects at all stages. Note that the parameters are estimated separately using estimating equations recursively at each stage. Therefore, it is natural to establish the asymptotic properties of the stage-specific parameters recursively as well. For simplicity, we will focus on the two-stage setting, but extensions to the general $T$-stage setting follow directly.
Theorem 1
In the 2-stage scenario, if Assumptions 1-4 hold, the Q-functions at stage 2 are Lipschitz continuous and correctly specified, and $\bm{M}_{2}(\bm{\theta}_{2})=\bm{0}$ has a unique solution, then
$\hat{\bm{\theta}}_{2}-\bm{\theta}_{2}^{*}$ $\stackrel{{\scriptstyle d}}{{\longrightarrow}}N(\bm{0},\Sigma_{\bm{\theta}_{2}}),$
$\hat{d}_{2}^{\mathrm{opt}}(\bm{h}_{2})$ $\stackrel{{\scriptstyle p}}{{\longrightarrow}}d_{2}^{\mathrm{opt}}(\bm{h}_{2})\text{ for all }\bm{h}_{2}\in\mathcal{H}_{2},$
$\hat{Y}_{pse,1}$ $\stackrel{{\scriptstyle p}}{{\longrightarrow}}Y_{pse,1},$
where $\bm{\theta}_{2}^{*}$ stands or the true values of $\bm{\theta}_{2}$. $\hat{\bm{\theta}}_{2}$, $\hat{d}_{2}^{\mathrm{opt}}(\cdot)$, and $\hat{Y}_{pse,1}$ denote the estimators of $\bm{\theta}_{2}$, $d_{2}^{\mathrm{opt}}(\cdot)$, and $Y_{pse,1}$, respectively. Notations $\stackrel{{\scriptstyle d}}{{\longrightarrow}}$ and $\stackrel{{\scriptstyle p}}{{\longrightarrow}}$ stand for converge in distribution and converge in probability, respectively. The detailed form of $\Sigma_{\bm{\theta}_{2}}$ is given in Web Appendix A.
Theorem 2
In the two-stage scenario, if Assumptions 1-5 hold, all models are Lipschitz continuous and correctly specified, and the estimating equations for the Q-functions and the working model (1) each have a unique solution, then under certain additional regularity conditions,
$\hat{\bm{\theta}}_{1,EE}-\bm{\theta}_{1}^{*}$ $\stackrel{{\scriptstyle d}}{{\longrightarrow}}N(\bm{0},\Sigma_{\bm{\theta}_{1}}),$
$\hat{d}_{1,EE}^{\mathrm{opt}}(\bm{h}_{1})$ $\stackrel{{\scriptstyle p}}{{\longrightarrow}}d_{1}^{\mathrm{opt}}(\bm{h}_{1})\text{ for all }\bm{h}_{1}\in\mathcal{H}_{1},$
where $\bm{\theta}_{1}^{*}$ stands or the true values of $\bm{\theta}_{1}$. $\hat{\bm{\theta}}_{1,EE}$ and $\hat{d}_{1,EE}^{\mathrm{opt}}(\cdot)$ denote the estimators of $\bm{\theta}_{1}$ and $d_{1}^{\mathrm{opt}}(\cdot)$ with the working model (1) estimated by the EE method, respectively. Notations $\stackrel{{\scriptstyle d}}{{\longrightarrow}}$ and $\stackrel{{\scriptstyle p}}{{\longrightarrow}}$ stand for converge in distribution and converge in probability, respectively. The regularity conditions and the detailed form of $\Sigma_{\bm{\theta}_{1}}$ are given in Web Appendix B.
Theorem 3
In the two-stage scenario, if Assumptions 1-4 and 6 hold, all models are Lipschitz continuous and correctly specified, the sensitivity parameter $\gamma^{\prime}$ is correctly specified, and the estimating equations for the Q-functions each have a unique solution, then under certain additional regularity conditions,
$\hat{\bm{\theta}}_{1,SA}-\bm{\theta}_{1}^{*}$ $\stackrel{{\scriptstyle d}}{{\longrightarrow}}N(\bm{0},\Sigma^{\prime}_{\bm{\theta}_{1}}),$
$\hat{d}_{1,SA}^{\mathrm{opt}}(\bm{h}_{1})$ $\stackrel{{\scriptstyle p}}{{\longrightarrow}}d_{1}^{\mathrm{opt}}(\bm{h}_{1})\text{ for all }\bm{h}_{1}\in\mathcal{H}_{1},$
where $\bm{\theta}_{1}^{*}$ stands or the true values of $\bm{\theta}_{1}$. $\hat{\bm{\theta}}_{1,SA}$ and $\hat{d}_{1,SA}^{\mathrm{opt}}(\cdot)$ denote the estimators of $\bm{\theta}_{1}$ and $d_{1}^{\mathrm{opt}}(\cdot)$ with the working model (4) estimated by the SA method, respectively. Notations $\stackrel{{\scriptstyle d}}{{\longrightarrow}}$ and $\stackrel{{\scriptstyle p}}{{\longrightarrow}}$ stand for converge in distribution and converge in probability, respectively. The regularity conditions and the detailed form of $\Sigma^{\prime}_{\bm{\theta}_{1}}$ are given in Web Appendix C.
Note that the form of the asymptotic variance of the parameters at stage 1 is complicated. Therefore, we suggest using the bootstrap method to estimate it in practice.
3 Simulation
We conducted two simulation studies to evaluate the finite-sample performance of the WQ-EE method. In Simulation 1, we focus on the scenario where assumptions 1-5 are satisfied. In Simulation 2, we evaluate the robustness of the WQ-EE method against the violations of Assumption 5 when (1) the chosen instrumental variable is weakly associated with the missingness probability and (2) there are interactions between $\bm{u}_{t}$ and $y_{pse,t}$ in the true missingness probability model.
For comparison, we also assessed three other methods for handling partially missing covariate data in Q-learning: a naive method which ignores partially observed covariates (‘naive’), complete case analysis (CC), and multiple imputation (MI) (Little and Rubin, 2014). Q-learning when all the covariates are fully observed (All) is used as the benchmark. In the naive method, we omitted all terms containing the partially observed covariates in the specified Q-functions. Consequently, the Q-functions are misspecified in this scenario as Assumption 2 is violated. In the CC method, we excluded all patients whose covariates are subject to missingness in estimation. In the MI method, we first imputed missing covariates using predictive mean matching multiple times, which was implemented through the R package MICE. Subsequently, we applied the standard Q-learning algorithm to the imputed data sets, and finally, we calculated the average of the estimated parameters from the imputed datasets to derive the final estimates of the Q-function parameters.
All methods were evaluated in two aspects: Firstly, we compared the performances of the methods for estimating the blip function parameters $\bm{\psi}_{t}$. Secondly, we compared their abilities to identify the true optimal DTR and the values of the estimated optimal DTRs.
3.1 Simulation 1: comparative performance when Assumptions 1-5 are satisfied
We simulated data from an observational study with two stages of intervention. There are two covariates $X_{t,1}$ and $X_{t,2}$ at stage $t=1,2$. Among these covariates, only $X_{1,2}$ and $X_{2,2}$ are partially observed, and their missingness indicators are $R_{1}$ and $R_{2}$, respectively. Denote $\text{expit}(b)=\exp(b)/(1+\exp(b))$ for $b$ in $\mathcal{R}$. The variables are generated as follows:
$\left(\begin{matrix}X_{1,1}\\ X_{2,1}\end{matrix}\right)\sim N\left(\left(\begin{matrix}0\\ 0\end{matrix}\right),\left(\begin{matrix}1&0.5\\ 0.5&1\end{matrix}\right)\right);$
$X_{1,2}\sim\text{Uniform}(0,2);R_{1}\sim\text{Bernoulli}([1+\exp(-3+X_{1,2})]^{-1});$
$A_{1}\sim 2\times\text{Bernoulli}(\text{expit}(-1+X_{1,1}+X_{1,2}-R_{1}))-1;$
$Y_{1}=A_{1}(\psi_{10}+2+\psi_{11}X_{1,2})+\beta_{10}+\beta_{11}X_{1,1}+\beta_{12}X_{1,2}+\epsilon_{1};\epsilon_{1}\sim N(0,3);$
$X_{2,2}\sim\text{Uniform}(0,2);R_{2}\sim\text{Bernoulli}([1+\exp(-1+0.5X_{1,2}-Y_{1}-X_{2,1}-0.5X_{2,2})]^{-1});$
$A_{2}\sim 2\times\text{Bernoulli}(\text{expit}(-1-X_{1,1}-X_{1,2}+Y_{1}+X_{2,1}-R_{2}))-1;$
$Y_{2}=A_{2}(\psi_{20}+\psi_{2,A}A_{1}+\psi_{22}X_{2,2})+\beta_{2,A}A_{1}+\beta_{22}X_{2,1}+\beta_{23}X_{2,2}+\epsilon_{2};\epsilon_{2}\sim N(0,1).$
The true values of the parameters $\bm{\theta}_{1}=(\psi_{10},\psi_{11},\beta_{10},\beta_{11},\beta_{12})$ and $\bm{\theta}_{2}=(\psi_{20},\psi_{2,A},\psi_{22},\beta_{2,A},\beta_{22},\beta_{23})$ were $(1,-1,1.5,0.5,-0.5)$ and $(1,-1,1,-1,1,-0.5)$, respectively. We generated $1000$ data sets with sample sizes of $500$ and $2000$.
We consider $Y_{pse,2}=Y_{1}+Y_{2}$ as the final outcome. Under this setting, the true optimal treatment rule at stage 2 is $12\mathbb{I}(\psi_{20}+\psi_{2,A}a_{1}+\psi_{22}x_{2,2}\geq 0)-$ 1. Because $\bm{\theta}_{2}=(1,-1,1,-1,1,-0.5)$ and $1-a_{1}+x_{2,2}\geq 0$, the optimal choice of $A_{2}$ is always 1. Then, the pseudo-outcome at stage $1$, $Y_{pse,1}$ , is equal to $1-2A_{1}+Y_{1}+X_{2,1}+0.5X_{2,2}$. Therefore, $Q_{1}(x_{1,1},x_{1,2},a_{1})=1.5+\beta_{10}+a_{1}(\psi_{10}+\psi_{11}x_{1,2})+(0.5+\beta_{11})x_{1,1}-\beta_{1,2}x_{1,2}$. And the optimal treatment rule at stage 1 is $12\mathbb{I}(\psi_{10}+\psi_{11}X_{1,2}\geq 0)-$ 1. Note that the missingness indicators of $X_{1,2}$ and $X_{2,2}$ are directly associated with their own values, thus they are missing not at random. $R_{1}$ only depends on $X_{1,2}$ and $R_{2}$ only depends on $(X_{1,2},Y_{1},X_{2,1},X_{2,2})$ so the future-independent missingness assumption holds. The missingness indicator for $Y_{pse,1}$ is the same as that for $X_{2,2}$, since $Y_{pse,1}=1-2A_{1}+Y_{1}+X_{2,1}+0.5X_{2,2}$ and $(A_{1},Y_{1},X_{2,1})$ are fully observed. Besides, since $P(R_{2}=1\mid\bm{H}_{2})=\{1+\exp(-1+0.5X_{1,2}-Y_{1}-X_{2,1}-0.5X_{2,2})\}^{-1}$, it follows that $P(R_{pse,1}=1\mid\bm{H}_{1},A_{1})=\left\{1+\exp(0.5X_{12}-2A_{1}-Y_{pse,1})\right\}^{-1}$, which indicates that $Y_{pse,1}$ is MNAR and $X_{1,1}$ is conditional independent of $R_{pse,1}$ given $(X_{1,2},A_{1},Y_{pse,1})$. Since $X_{1,1}$ is also associated with $Y_{pse,1}$ after conditioning on $(X_{1,2},A_{1},Y_{pse,1})$, we employ $X_{1,1}$ as a nonresponse instrument variable.
Figure 3: Boxplots of blip function parameter estimates from the compared methods in Simulation 1. In each boxplot, white and grey boxes are the results for the sample sizes 500 and 2,000, respectively. The horizontal line marks the true values of the blip function parameter estimates. Note that the naive method did not generate estimates of $\psi_{11}$ and $\psi_{22}$ as $\psi_{11}$ and $\psi_{22}$ are the coefficients of partially missing $X_{1,2}$ and $X_{2,2}$ in the true treatment rules.
Figure 3 shows the empirical distributions of the estimates of $\bm{\psi}_{t}$, which directly influence the estimated optimal DTR. At stage 2, the proposed WQ-EE method and the CC analysis yielded unbiased estimates of the blip function parameters since the future-independent missingness assumption holds and the outcome at stage 2 is fully observed. In contrast, the naive and MI methods led to non-negligible biases of the blip function parameters. At stage 1, the estimators in the WQ-EE method remained unbiased while the estimators in the naive, CC, and MI methods displayed relatively large biases, which were not mitigated as the sample size increases. Note that the naive estimates of $\psi_{11}$ and $\psi_{22}$ are unavailable, as we excluded the terms $A_{1}X_{1,2}$ and $A_{2}X_{2,2}$ from the specified Q-functions when implementing this method.
The proportion of patients whose optimal treatments in both stages were correctly identified was high when employing the WQ-EE method, as demonstrated in Table 1. With the smaller sample size ($n=500$), the WQ-EE method achieved an average correct classification rate of $89.9\%$ across the 1000 simulated datasets, with a mean final outcome of $2.98$. In general, larger sample sizes led to higher proportions of patients with their true optimal DTR correctly identified. In contrast, the DTRs obtained from CC and MI methods had much lower correct classification rates and the mean final outcomes under the estimated DTRs were lower than that from the WQ-EE method.
Table 1: Performance of estimated DTRs in Simulation One
n=500 n=2000
Value OptValue Opt%
All 2.988 (0.057) 0.928 (0.052) 2.997 (0.028) 0.964 (0.024)
naive 2.501 (0.053) 0.500 (0.022) 2.500 (0.027) 0.499 (0.011)
CC 2.879 (0.128) 0.773 (0.111) 2.898 (0.059) 0.777 (0.057)
MI 2.832 (0.139) 0.729 (0.108) 2.849 (0.071) 0.730 (0.058)
WQ-EE 2.974 (0.069) 0.899 (0.073) 2.994 (0.029) 0.949 (0.036)
  • Note: Value, the value of the estimated regime; Opt%, the percentage of patients correctly classified to their true optimal treatments in both stages; Std, the sample standard error.
3.2 Simulation 2: comparative performance when Assumption 5 is violated
To investigate the robustness of the proposed WQ-EE method when Assumption 5 is violated, we kept all other data generating mechanisms unchanged, while altering the true model for $R_{pse,1}$ to $P(R_{pse,1}=1)=\left\{1+\exp(\gamma_{z}X_{1,1}+0.5X_{1,2}-2A_{1}-Y_{pse,1})\right\}^{-1}$. When $\gamma_{z}=0$, the data generating mechanism is the same as that in Simulation 1 and Assumption 5 is satisfied. However, when $\gamma_{z}\neq 0$, $X_{1,1}$ is associated with $P(R_{pse,1}=1)$ conditional on $(X_{1,2},A_{1},Y_{pse,1})$. Thus there are no nonresponse instrument variables. We set $\gamma_{z}\in\{-0.4,-0.2,0.2,0.4\}$ such that $X_{1,1}$ is weakly associated with $P(R_{pse,1}=1)$ conditional on on $X_{1,2}$ and $Y_{pse,1}$. We generated 1000 Monte Carlo samples with sample size $n=500$.
Since the future-independent missingness assumption still holds in Simulation 2, the estimators for $(\psi_{20},\psi_{2,A},\psi_{22})$ in the WQ-EE method are still consistent. We report the performance of the estimators for $(\psi_{10},\psi_{11})$ and the value and correct classification rate of the estimated optimal DTRs. The results were calculated by averaging over 1000 simulated data sets and are shown in Table 1 2. We see that the proposed WQ-EE method was robust when $X_{11}$ was weakly related to $R_{pse,1}$. Despite the declining performance of the corresponding optimal DTR as the absolute value of $\gamma_{z}$ increased, the WQ-EE method consistently outperformed the other 3 methods with higher mean final outcomes and correct classification rates.
Table 2: Performance of the estimators without nonresponse instrumental variable based on 1000 Monte Carlo samples in Simulation Two
$\psi_{10}$ $\psi_{11}$ Value Opt%
Method Bias Std Bias Std Mean Std Mean Std
$\gamma_{z}=-0.4$
naive -1.005 0.168 2.498 0.054 0.499 0.022
CC 0.068 0.188 -0.447 0.277 2.881 0.131 0.773 0.113
MI 0.166 0.169 -0.588 0.242 2.838 0.144 0.738 0.116
WQ-EE -0.042 0.237 0.113 0.318 2.971 0.078 0.896 0.080
$\gamma_{z}=-0.2$
naive -1.008 0.166 2.498 0.054 0.500 0.023
CC 0.062 0.184 -0.443 0.266 2.885 0.126 0.781 0.113
MI 0.161 0.165 -0.591 0.231 2.838 0.142 0.733 0.113
WQ-EE -0.043 0.243 0.090 0.315 2.972 0.081 0.898 0.076
$\gamma_{z}=0.2$
naive -1.000 0.155 2.502 0.052 0.499 0.022
CC 0.073 0.185 -0.482 0.253 2.871 0.125 0.766 0.110
MI 0.174 0.164 -0.627 0.227 2.816 0.144 0.718 0.113
WQ-EE -0.040 0.239 0.031 0.310 2.974 0.070 0.899 0.072
$\gamma_{z}=0.4$
naive -1.007 0.151 2.501 0.053 0.499 0.023
CC 0.088 0.185 -0.495 0.274 2.862 0.132 0.755 0.114
MI 0.183 0.163 -0.639 0.238 2.808 0.146 0.705 0.114
WQ-EE -0.020 0.233 0.053 0.321 2.968 0.074 0.895 0.078
  • Note: Value, the value of the estimated regime; Opt%, the percentage of test subjects correctly classified to their true optimal treatments in both stages; Std, the sample standard error.
We also consider scenarios where there are interactions between $y_{pse,t}$ and $\bm{u}_{t}$ in the true missingness probability model. The true model for $R_{pse,1}$ was modified as $P(R_{pse,1}=1)=[1+\exp(0.5X_{1,2}-2A_{1}-Y_{pse,1}+\gamma_{uy}(X_{1,2}-1)Y_{pse,1})]^{-1}$, while the data generating mechanism for other covariates was the same as that in Simulation 1. The specific form for the interaction, $(X_{12}-1)Y_{pse,1}$, was chosen to ensure that the missing data proportions were largely similar between Simulations 1 and 2. We set $\gamma_{uy}\in\{-0.2,-0.1,0.1,0.2\}$ by allowing weak interactions between $X_{1,2}$ and $Y_{pse,1}$. We again generated 1000 simulated data sets with sample size $n=500$. The results of the compared methods are shown in Table 3. The WQ-EE method exhibits superior performance compared to the other three methods despite the introduction of interaction terms involving $Y_{pse,t}$, showcasing its robustness in the presence of mild model specifications.
Table 3: Performance of the estimators when interaction terms between $y_{psr,t}$ and $\bm{u}_{t}$ are incorrectly omitted based on 1000 Monte Carlo samples in Simulation Two
$\psi_{10}$ $\psi_{11}$ Value Opt%
Method Bias Std Bias Std Est Std Est Std
$\gamma_{uy}=-0.2$
naive -1.011 0.160 2.500 0.053 0.500 0.022
CC 0.124 0.184 -0.514 0.269 2.863 0.139 0.760 0.115
MI 0.215 0.162 -0.656 0.233 2.808 0.152 0.712 0.115
WQ-EE -0.067 0.239 0.075 0.324 2.972 0.072 0.898 0.073
$\gamma_{uy}=-0.1$
naive -1.002 0.153 2.500 0.053 0.499 0.022
CC 0.084 0.182 -0.474 0.270 2.871 0.132 0.764 0.116
MI 0.185 0.159 -0.624 0.230 2.820 0.141 0.715 0.115
WQ-EE -0.062 0.228 0.065 0.318 2.971 0.072 0.897 0.075
$\gamma_{uy}=0.1$
naive -1.002 0.157 2.499 0.053 0.500 0.022
CC 0.046 0.182 -0.441 0.258 2.886 0.119 0.780 0.105
MI 0.147 0.163 -0.595 0.231 2.832 0.143 0.734 0.107
WQ-EE -0.021 0.233 0.022 0.307 2.974 0.074 0.898 0.077
$\gamma_{uy}=0.2$
naive -1.010 0.160 2.497 0.051 0.499 0.023
CC 0.046 0.188 -0.444 0.263 2.883 0.124 0.772 0.112
MI 0.146 0.168 -0.590 0.234 2.835 0.140 0.732 0.111
WQ-EE 0.023 0.242 -0.011 0.308 2.972 0.079 0.895 0.082
  • Note: Value, the value of the estimated regime; Opt%, the percentage of test subjects correctly classified to their true optimal treatments in both stages; Std, the sample standard error.
Overall, our simulation studies demonstrated that the proposed WQ-EE method performed well under correct model specification and mild model misspecification, in comparison with the naive, CC, and MI methods.
4 Application to the MIMIC-III data
We applied the proposed methods to the MIMIC-III data (Johnson et al., 2016) for investigating the optimal 2-stage fluid strategy to treat sepsis patients in ICUs.
Sepsis is a life-threatening clinical syndrome that occurs when the body’s response to an infection damages its own tissues and organs. Administering timely and efficient fluid resuscitation is crucial in stabilizing tissue hypoperfusion or septic shock induced by sepsis. The ‘Surviving Sepsis Campaign’, an established clinical guideline for sepsis treatment, strongly recommends the administration of at least 30 mL/kg of intravenous (IV) fluid during the first 3 hours of ICU admission (Rhodes et al., 2017). Motivated by the scarcity of strong evidence for the most beneficial fluid resuscitation strategy in the early hours of treatment, Speth et al. (2022) utilized MIMIC-III data to assess the optimal DTR for 0-3 hours and 3-24 hours post-admission to the medical ICU. Building upon Speth et al. (2022) and further aligning with the recommendation from the ‘Surviving Sepsis Campaign’ that additional fluid administration post-initial resuscitation be guided by reassessing the patient’s hemodynamic status, we included hemodynamic variables when estimating the optimal fluid resuscitation strategy during the 3-24 hours period post-ICU admission (i.e., stage 2).
We employed similar patient selection criteria as outlined in Speth et al. (2022) and focused on adult septic patients admitted to the medical ICU after initially presenting to the emergency department. The detailed cohort eligibility criteria can be found in Web Appendix E. The treatments in both stages, the baseline covariates, and the final outcome were defined according to Speth et al. (2022). Specifically, the stage 1 and stage 2 treatments were categorized as either fluid restrictive ($<$30 ml/kg) or fluid liberal ($\geq$ 30 ml/kg) strategies in 0-3 hours and 3-24 hours post-ICU admission, respectively. The baseline covariates included gender, age, weight, racial groups, and Elixhauser comorbidity score. The outcome of interest was the Sequential Organ Failure Assessment (SOFA) score evaluated at 24 hours post-admission. The SOFA score is a clinical tool used to assess the severity of organ dysfunction by assigning scores to various organ systems. A higher SOFA score indicates a more severe impairment. The value of the estimated regime was defined as the improvement of the SOFA score at 24 hours, which was the difference between the means of the observed and the expected SOFA scores at 24 hours post-admission under the optimal DTR.
For the intermediate variables considered before stage 2 treatment, we included mechanical ventilation and vasopressors within the first 3-hour period, as well as the patients’ SOFA score evaluated at 3 hours post admission. Additionally, we also considered heart rate, blood pressure, SpO2, respiratory rate, temperature at 3 hours post-admission and urine output within 0-3 hours after admission. These hemodynamic variables contained missing values. As discussed previously, these variables were likely to be MNAR due to informative monitoring of the patients in the MIMIC-III database. The missing proportion of each hemodynamic variable is provided in Web Table 3. There were 973 patients in the selected cohort, among them $53.2\%$ were male and $78.7\%$ were Caucasian. 45.5% of the patients received the fluid liberal strategy in 0-3 hours while $53.5\%$ received the fluid liberal strategy in 3-24 hours. Overall, 67.1% of patients had fully observed covariates at stage 2, while the rest of the patients had at least one covariate with missing value.
As discussed in Section 2.3, it is plausible that Assumption 4 holds, i.e., that the missing data mechanism of the medical history up to stage $t$ is unrelated to the patient’s medical records after stage $t$, conditioning on the patient’s medical history and treatments up to stage $t$. Denote the SOFA score at 24 hours under the optimal stage 2 treatment as $Y_{pse,1}$. It is plausible that weight and gender were independent of the missingness of hemodynamic variables given baseline covariates and $Y_{pse,1}$. Since weight and gender were associated with $Y_{pse,1}$, they were suitable candidates for nonresponse instrumental variables. Additionally, the Elixhauser comorbidity score can be regarded as a mismeasured proxy of $Y_{pse,1}$, which may also serve as a candidate for the nonresponse instrumental variable, according to Miao and Tchetgen (2018). We investigated the impact of using different sets of the candidates for nonresponse instrumental variables in Table 4. Besides, we applied CC, MI, and the naive method that ignored hemodynamic variables for comparison. The improvement of SOFA at 24 hours was then calculated as the difference between the means of the observed and the expected SOFA scores at 24 hours under the estimated optimal DTR. We performed 1000 bootstrap resamplings and applied the 5-fold cross-validation procedure to each bootstrap sample to assess the performance of the estimated optimal DTRs. More details of the estimation procedure are given in Web Appendix E.
Table 4: Improvement of SOFA at 24 hours for sepsis patients in the selected cohort of the MIMIC-III database under the estimated optimal fluid strategies with different approaches to handling MNAR covariates.
Mean Bootstrap std Bootstrap 95% CI
WQ-EE (ECS) 1.294 0.163 (0.971, 1.634)
WQ-EE (weight) 1.439 0.156 (1.141, 1.760)
WQ-EE (ECS, weight) 1.334 0.157 (1.032, 1.668)
WQ-EE (gender) 1.428 0.181 (1.113, 1.775)
WQ-EE (gender, ECS) 1.311 0.167 (1.052, 1.716)
WQ-EE (gender, weight) 1.462 0.165 (1.167, 1.799)
WQ-EE (gender, weight, ECS) 1.339 0.170 (1.044, 1.692)
CC 1.193 0.148 (0.924, 1.476)
MI 1.093 0.110 (0.884, 1.316)
naive 0.911 0.102 (0.785, 1.127)
  • Note: Mean, the mean of bootstrap point estimates under 5-fold cross-validation; Bootstrap Std, the sample standard error of 1000 cross-validation estimates with bootstrap samples; Bootstrap 95% CI, 95% confidence interval constructed by bootstrap percentiles; ECS, Elixhauser comorbidity score; WQ-EE ($\cdot$), the WQ-EE method based on the corresponding nonresponse instrumental variables listed in parenthesis.
Table 4 presents the bootstrap mean, standard error, and 95% confidence interval of the improvement of SOFA score at 24 hours. Notably, all three methods that considered hemodynamic variables outperformed the naive method, which aligns with the recommendation from the Surviving Sepsis Campaign. Furthermore, the estimates based on the WQ-EE method were notably larger than those based on the CC and MI methods. Importantly, while the WQ-EE method exhibited robustness to the choice of the three candidate nonresponse instrumental variables, the estimated improvement of SOFA at 24 hours would be lower if the Elixhauser comorbidity score was included in the set of nonrsponse instrumental variables. This could be attributed to the association between the Elixhauser comorbidity score and the missingness of hemodynamic variables. In Simulation 2, we have demonstrated that this association can affect the performance of the WQ-EE method. Therefore, we recommend excluding the Elixhauser comorbidity score from the set of nonresponse instrumental variables. The remaining three combinations of nonresponse instrumental variables showed similar means and standard errors for the improvement of SOFA at 24 hours.
However, we cannot rule out that gender and weight could be weakly associated with the missingness of hemodynamic variables, even after conditioning on other covariates and the pseudo-outcome. Therefore, we also applied the proposed WQ-SA method in case none of the aforementioned variables were valid nonresponse instrument variables. We assumed $\gamma^{\prime}_{1}\leq 0$ because patients with higher SOFA scores were more severely ill, and they may undergo more intensive monitoring of their hemodynamic status. We considered a range of values for $\gamma^{\prime}_{1}$ within the interval $[-2,0]$. The rationale for selecting this range was that when we included gender and weight as nonresponse instrumental variables, the estimated value of $\gamma_{1}$ in the working model (1) results in a bootstrap mean of $-0.630$ with a bootstrap standard error of $0.469$. We believe that the coefficient of $Y_{pse,1}$ would not change considerably if gender and weight were included in the working model (4) for the missingness probability. Furthermore, when $\gamma^{\prime}_{1}=-2$, it deviated by almost three times the standard error from the bootstrap mean, allowing a wide range for the residual association with $Y_{pse,1}$. Table 5 shows that when $\gamma^{\prime}_{1}=0$, the estimated improvement in SOFA at 24 hours was nearly identical to that of the CC method, which also assumed a future-independent missingness assumption and ignored the relationship between $Y_{pse,1}$ and $R_{pse,1}$. The estimated improvement in SOFA at 24 hours increased rapidly as $\gamma^{\prime}_{1}$ changed from 00 to $-1$. When $\gamma^{\prime}_{1}$ changes from -1 to -2, the estimated improvement in SOFA at 24 hours stabilized at around 1.5.
Table 5: The improvement of SOFA at 24 hours for sepsis patients in the selected cohort of the MIMIC-III database under the estimated optimal fluid strategies with different values of the sensitivity parameter $\gamma_{1}^{\prime}$ in the WQ-SA method
$\gamma^{\prime}_{1}$ Mean Bootstrap std Bootstrap 95% CI
0 1.185 0.151 (0.894, 1.479)
-0.5 1.349 0.160 (1.060, 1.680)
-1 1.461 0.177 (1.130, 1.827)
-1.5 1.511 0.185 (1.172, 1.891)
-2 1.559 0.209 (1.198, 2.027)
  • Note: Mean, the mean of bootstrap point estimates under 5-fold cross-validation; Bootstrap Std, the sample standard error of 1000 cross-validation estimates with bootstrap samples; Bootstrap 95% CI, 95% confidence interval constructed by bootstrap percentiles.
While the challenge of optimizing the management of septic patients is intricate and multifaceted, we employed the proposed methods to investigate whether incorporating partially missing hemodynamic variables could influence the fluid administration after the initial resuscitation to enhance the overall outcome of the septic patients. In line with the best practice recommendations of the Surviving Sepsis Campaign, our analysis results indicated that certain hemodynamic variables should be taken into account when formulating the fluid resuscitation strategy during the 3-24 hours following admission to the medical ICU.
5 Discussion
In this article, we proposed two weighted Q-learning approaches to estimating optimal DTRs with MNAR covariates and consequent MNAR pseudo-outcomes. To the best of our knowledge, this is the first attempt to estimate optimal DTRs under these challenging MNAR conditions. Specifically, we utilized the future-independent missingness assumption and developed a semi-parametric estimating equation-based approach to estimating the conditional missingness probability of the pseudo-outcomes with valid nonresponse instruments. After obtaining the estimated missingness probabilities, we applied inverse probability weighting to estimate the Q-function parameters. When nonresponse instrumental variables were not available, we specified a sensitivity parameter in the working model for missingness probability and conducted a sensitivity analysis to assess the robustness of the estimated DTRs. In our simulation studies, we focused on evaluating the performance of the WQ-EE method as the performance of the WQ-SA method would rely on the correct specification of the sensitivity parameter. Our simulation results showed that the estimators in the WQ-EE method were unbiased when Assumptions 1-5 hold. They were also robust to mild model misspecification when Assumption 5 was violated, while the alternative methods led to sub-optimal DTRs when covariates were MNAR. In the application to the MIMIC-III data, we showed that incorporating hemodynamic variables into optimal fluid strategy estimation and handling MNAR problems with the proposed methods can improve the SOFA score for sepsis patients in ICUs.
There are several directions for future work. First, this article exclusively addresses the issue of missing pseudo-outcomes due to MNAR covariates in DTR estimation with Q-learning. However, in observational studies or EMR databases, patients may not progress through all stages of clinical intervention due to experiencing a terminal event or being censored before the end of the study. In such cases, it is desirable to integrate the terminal event and longitudinal outcome measurements for estimating DTRs while accounting for the censoring and MNAR covariates. Second, the dimension of the covariates is often high in practice, especially in multiple-stage settings. In such scenarios, the kernel regression method is susceptible to the ‘curse of dimensionality’ problem. To address this challenge, one may consider dimension reduction techniques proposed in Tang et al. (2014). Third, the approach proposed for handling MNAR covariates and pseudo-outcomes can be extended to other dynamic programming methods for DTRs, such as A-learning.